OpenStudy (anonymous):

Determine whether the vectors u and v are parallel , orthogonal , or neither.

OpenStudy (anonymous):

u=(1,2) v=(-4,8)

OpenStudy (anonymous):

I know the answer is not orthogonal because the dot product does not = 0. I need help in determining if it's parallel

OpenStudy (jhannybean):

Parallel vectors mean the angle between them is \(0^\circ\) Orthogonal (perpendicular) vectors have an angle between them that = \(90^\circ\).

OpenStudy (anonymous):

OK I had it mixed up then. So they are not parallel

OpenStudy (jhannybean):

So in order to determine this, you need to use the dot product and the cross product of vectors that give you an angle measurement. Dot Product: \(\vec a \cdot \vec b = |a||b|\cos(\theta)\) Cross product: \(\vec a \times \vec b = |a||b|\sin(\theta)\)

OpenStudy (anonymous):

So u = -4, and v= -16

OpenStudy (jhannybean):

\[\theta = \cos^{-1}\left(\frac{\vec a \cdot \vec b}{|\vec a||\vec b|}\right) \]\[\theta = \sin^{-1}\left(\frac{\vec a \times \vec b}{|\vec a||\vec b|}\right)\]

OpenStudy (jhannybean):

and theres a typo in my first post. It should be \[\vec a \cdot \vec b = |\vec a||\vec b|\cos(\theta)\]\[\vec a \times \vec b = |\vec a||\vec b|\sin(\theta)\]

OpenStudy (jhannybean):

So i'll start you off. \[\vec u = \langle a_1~,~a_2\rangle \implies \vec u = \langle 1~,~2 \rangle\]\[\vec v = \langle b_1~,~b_2 \rangle \implies \vec v =\langle -4~,~8\rangle\] \[\vec u \cdot \vec v = (a_1b_1) +(a_2b_2)\]\[|\vec u| = \sqrt{a_1^2+a_2^2}\]\[|\vec v| = \sqrt{b_1^2+b_2^2}\]

OpenStudy (jhannybean):

now just solve for all these parts and plug them in to the formula. Depending on whether you get \(0^\circ\) or \(90^\circ \) , you can determine whether its parallel or perpendicular.

OpenStudy (anonymous):

The 2 is suppose to be a -2 that was a typo -4+(-16)= -20

OpenStudy (jhannybean):

Which part are you solving for?

OpenStudy (anonymous):

|u| = 2.23 or sqrt 5 |v|= 8.94 or 4 sqrt 5

OpenStudy (jhannybean):

\[|\vec u| = \sqrt{1^2+(-2)^2} = \sqrt{5} =2.2~\checkmark \]\[|\vec v| =\sqrt{(-4)^2+8^2} = \sqrt{16+64}= \sqrt{80}=4\sqrt{5} =8.94 ~\checkmark\]

OpenStudy (jhannybean):

Next step?

OpenStudy (jhannybean):

Now you should find \[\vec u \cdot \vec v\]\[\vec u \times \vec v\]

OpenStudy (anonymous):

I'm not sure. Do I do something with the -20 answer I got previously?

OpenStudy (anonymous):

2.2 × 8.94= 19.66

OpenStudy (jhannybean):

Oh!!! I forgot something major. :(

OpenStudy (jhannybean):

It's the MAGNITUDE of the cross product! \[|\vec u \times \vec v| = |\vec u||\vec v|\sin(\theta)\]

OpenStudy (jhannybean):

And so you can test whether the angle between the perpendicular vectors is orthogonal by taking the magnitude of their cross product. \[|\vec u \times \vec v| = 0 \implies \vec u~ ||~\vec v\]

OpenStudy (anonymous):

Now I'm confuse...

OpenStudy (jhannybean):

You've found \(|\vec u|\) and \(|\vec v|\), now find \[\vec u \cdot \vec v\]\[|\vec u \times\vec v|\]

OpenStudy (anonymous):

19.66?

OpenStudy (jhannybean):

And what exactly is 19.66....?

OpenStudy (anonymous):

What do you mean?

OpenStudy (jhannybean):

\(\color{blue}{\text{Originally Posted by}}\) @allydiaz 19.66? \(\color{blue}{\text{End of Quote}}\) What are you referring to?

OpenStudy (anonymous):

Sorry your reply looks weird on my computer

OpenStudy (jhannybean):

refresh maybe?

OpenStudy (anonymous):

Is there something I do next after 19.66. Or does that mean it's orthogonal ?

OpenStudy (jhannybean):

How did you get 19.66? It looks like a random number. What part did you solve to get 19.66? Thats what I am asking you to clarify.

OpenStudy (irishboy123):

u=(1,2), |u|= root(5) v=(-4,8), |v| = root (80) u . v = |u| |v| cos t cos t = {u . v }/ {|u| |v|} u . v = -4 + 16 = 12 cos t = 12 / { root(400) } = 3/5 t = ... doesn't look like they are parallel or ortogonal..

OpenStudy (jhannybean):

Yep, that's what I got as well, I was looking for the OP to solve the equation.

OpenStudy (anonymous):

U × v= 2.23 × 8.94

OpenStudy (anonymous):

The 2 is suppose to be a negative 2. I stated stated that i was typo in an earlier post

OpenStudy (anonymous):

@Jhannybean and @IrishBoy123

OpenStudy (anonymous):

So it doesn' t matter that its u=(1,-2) ?

OpenStudy (irishboy123):

"So in order to determine this, you need to use the dot product and the cross product of vectors that give you an angle measurement. " pasting formulae off the internet is not going to help anyone. @allydiaz look at my last post. you only need to use the scalar/dot product.

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

I still get my same answer : (

OpenStudy (irishboy123):

of course it matters. massively. u = <1, -2>; |u| = root(5) v = < -4, 8>, |v| = root ( 80) u.v = -4 - 16 = -20 = root(400) cos t = 20 cost cost = -1 t = - 180 deg. pointing in different directions. these things are parallel!! all you need to understand is the dot product to get that.

OpenStudy (anonymous):

Thank you!! That's what I got as well

OpenStudy (irishboy123):

cool!